# An Empirical Study on Compliance with Ranking Transparency in the Software Documentation of EU Online Platforms

### 22 Dec 2023 v1

@@ Abstract.
Compliance with the European Union’s Platform-to-Business (P2B) Regulation is challenging for online platforms, and assessing their compliance can be difficult for public authorities. This is partly due to the lack of automated tools for assessing the information (e.g., software documentation) platforms provide concerning ranking transparency. Our study tackles this issue in two ways. First, we empirically evaluate the compliance of six major platforms (Amazon, Bing, Booking, Google, Tripadvisor, and Yahoo), revealing substantial differences in their documentation. Second, we introduce and test automated compliance assessment tools based on ChatGPT and information retrieval technology. These tools are evaluated against human judgments, showing promising results as reliable proxies for compliance assessments. Our findings could help enhance regulatory compliance and align with the United Nations Sustainable Development Goal 10.3, which seeks to reduce inequality, including business disparities, on these platforms.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: 46th International Conference on Software Engineering; April 2024; Lisbon, Portugalprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Documentationccs: Information systems Data analyticsccs: Applied computing Lawccs: Computing methodologies Natural language processing
@@
In today’s digital world, software powers online marketplaces, search engines, and platforms businesses use daily. When you search for a product on Amazon or a hotel on Booking, the software determines which options show up first. It is vital for businesses to understand these systems, but the explanations are often unclear. The European Union mandated that online platforms explain how their ranking systems work. This promotes fairness and reduces inequalities. Yet, ensuring platforms adhere can be challenging due to the absence of a standard assessment method. We investigated how giants like Google, Yahoo, Bing, Amazon, Booking, and Tripadvisor clarify their ranking systems. We asked three legal experts and more than 100 people to evaluate these explanations. The result? The clarity varies significantly between platforms. Manually checking this is laborious. So, we developed tools to help. We introduced two automated tools and tested their accuracy against human evaluations, discussing how they can enhance online platform transparency. Clearer explanations lead to a fairer digital environment for businesses, supporting global goals of reducing inequality.
@@ 1. Introduction
To address this issue, the European Union has introduced legislation such as the Platform-to-Business (P2B) Regulation (EU) 2019/1150 (European Parliament and Council, 2019). This Regulation requires platforms to provide easily accessible, comprehensible, and detailed documentation explaining how ranking mechanisms work and affect sellers, including business and corporate website users ( Article 5 and Recital 25 (Commission, 2023)). Such a measure aims to create a more equitable environment for businesses that rely on online platforms to connect with consumers. This resonates with Target 10.3 of the United Nations Sustainable Development Goals (of Economic and Affairs, 2023), which aims to ensure equal opportunity and reduce inequalities within and among countries.
However, producing documentation adhering to the P2B Regulation is challenging for online platforms, as recently pointed out by the European Commission (EC) (Commission, 2023). Additionally, the quality of such documentation is difficult to monitor, both for the platforms themselves and for regulators. This is largely due to the lack of standardized, automated tools for assessing compliance across different platform providers. Hence, given the absence of standardized assessment tools, we hypothesize to encounter varying levels of compliance among platforms regarding ranking transparency.
@@
    Source code for the automated tools and referenced data, including manual assessments and checklist (Sovrano, 2023) [^1] https://github.com/Francesco-Sovrano/Automating-Regulatory-Compliance-An-Empirical-Study-on-Ranking-Transparency-of-EU-Online-Platforms.
@@
Structure of the paper. Section 2 discusses related work in software documentation, legal compliance, and automated assessment tools. Section 3 offers background on the P2B Regulation and the technology for our assessment tools. Section 4 describes our study design and the checklist for documentation assessment. Section 5 introduces two algorithms for automated assessment of software documentation. In Section 6, we share the outcomes of both manual and automated evaluations. Section 7 considers the implications of our findings and threats to validity, with societal impacts highlighted in Section 8. Lastly, Section 9 concludes and suggests future research directions.
@@ 2. Related Work
Additionally, Lukovic (Lukovic, 2021) examined the context and actors of platform rankings. This study details the ranking transparency obligations set by the P2B Regulation and offers critical insights into the framework’s limitations. Bibal et al. (2021) situate the P2B Regulation’s ranking parameters within a broader EU context that mandates explanations for AI systems. This work sheds light on the intentions of EU legislators and specifies the type and extent of information essential for technical compliance. The approaches of both Lukovic (2021) and Bibal et al. (2021) differ from ours as they neither evaluate platforms’ adherence to the Regulation nor introduce tools to enhance regulatory compliance.
@@ 3. Background
This section covers the P2B Regulation and the DoX metric. We delve into the P2B Regulation, an EU law promoting fairness between online platforms and businesses, and touch on the European Commission’s ranking transparency guidelines. Then, we discuss the DoX metric, which measures explainability. This background prepares readers for the assessment tools discussed later.
3.1. P2B Regulation & European Commission Guidelines on Ranking Transparency
@@
The EU legislator aims to ensure that online platforms provide explanations, enabling business and corporate website users to grasp key aspects of ranking mechanisms (Commission, 2023; Shaping Europe’s digital future, 2023). These include understanding how ranking considers the offers’ characteristics, their relevance to consumers, and, for search engines, website design characteristics (Art. 5.5). The explanations must also highlight the potential to influence ranking via payments and their effects (Art. 5.3).
@@
Regarding the form that the explanations should take, both the Regulation and the Guidelines use the term ’drafted’ in English, which refers to written content. [^2] See, for instance, the definitions and examples proposed at https://dictionary.cambridge.org/fr/dictionnaire/anglais/draft, about the verbal (and especially transitive) form of the word draft. To verify this, we looked into the French versions of both texts. The French translation of the Regulation refers to redacted explanations (‘rédigées’) within their terms and conditions for online intermediation services (Art. 3.1 and 5.1). For search engines, the exact text uses the words ‘to state’ (‘énoncée’), which do not imply the use of one form over another (Art. 5.2). The Guidelines, however, refer at some point to the ‘redaction’ of the explanations for both types of providers, potentially implying the use of a written medium in all cases (i.e., “lorsqu’ils rédigent la description […]”, pt. 99). We also looked into the Dutch version of the Regulation, which refers to ‘redacted’ content (‘opgesteld’) regardless of the type of service provider (Art. 3.1 and 5.2).
@@ 3.2. How to Measure Explainability
To calculate DoX scores, the relevance of a text snippet is assessed and aggregated based on its relation to a collection of archetypal questions informed by linguistic theories (see (Sovrano and Vitali, 2023)). In detail, these relevance measurements are obtained employing pre-trained deep language models tailored for general-purpose answer retrieval (Karpukhin et al., 2020; Bowman et al., 2015).
@@ 4.1. The Study
Our selection of the six platforms is driven by representativeness and audience profile. For representativeness, we target key players under the P2B Regulation, encompassing both intermediation services and search engines. Three platforms focus on intermediation, the others on search, all influential within the EU. Regarding audience profile, we assess how platforms explain ranking mechanisms to their users, emphasizing business visitors. Following pt. 17 of the Guidelines, explanations should cater to the ‘average’ user, balancing clarity and depth. All selected platforms serve a non-specialist audience.
@@ 4.2. The Checklist
To evaluate the quality of the documentation of the selected platforms, we created a checklist of essential questions, which should all be answered positively to conclude that an operator complies with ranking transparency requirements. These questions were designed by a legal expert, co-author of this study, based on the objectives and wording of the obligations imposed by the P2B regulation and the guidelines issued by the European Commission (Section 3.1). To elaborate on them, we adopted the following methodology.
@@
However, as pointed out in Section 3.1, the Regulation differentiates between two types of platform operators: online intermediation services and online search engines. While both are covered by the Regulation, they face slightly different requirements, as elaborated in Articles 5.1 and 5.2, and in point 7 of the Guidelines. Therefore, we have created a different set of yes-or-no questions for each platform type, as detailed below. Importantly, this task was led by solely paying attention to the explicit requirements contained in the Regulation.
Thirdly, we drafted additional questions based on the content of the European Commission’s guidelines. Whenever these guidelines explicitly raised new questions, proposed nuances to the requirements of the legal text, or gave best practices to comply with the P2B Regulation and its objectives, we designed new questions to take these elements into account and incorporated them into our checklist. These questions relate to elements like (i) the explanation of what is most important in determining ranking (cf. Q12) and (ii) the internal process conducted by the platforms to determine the main parameters (cf. Q14).
@@ 2. Q19
Although some of the questions, within each subset, might seem quite similar at first sight, they all call for (at least slightly) different answers, as they contain nuances that originate from the European Commission’s Guidelines. For instance, questions number 4 and 13 appear to be very similar. However, question 4 puts emphasis on the contrast between the parameters selected as main ones and other parameters, whereas question 13 does not. Similarly, questions number 3 and 11 seem quite close to each other. Yet, question 3 only refers to the main parameters used, while question 11 additionally considers how the ranking mechanism works in general.
@@ 5.2. Assessing Explanation Quality with DoX
In particular, we employ the same information retrieval system used by the DoX algorithm (cf. Section 3.2) to extract the answers to the checklist questions from the documents and ChatGPT to refine them into comprehensive explanations. Then, DoX is used to measure the explanatory quality of the answers to the checklist questions. Using the earlier example for clarity: while both answers ‘Rankings are based on algorithms’ and ‘Rankings are determined by algorithms that consider factors like user engagement, relevance, and content quality’ would achieve similar relevance ratings, the first would score lower on the DoX metric, indicating its limited explanatory depth.
@@
Paragraphs: {contents} This template is designed for re-elaborating the output of an automated answer retrieval system like the one we employed. The prompt consists of guidelines on how to form a comprehensive answer based on the paragraphs provided for a specific question. It asks the system to generate an answer solely based on the information found in these paragraphs. It also requests that the source paragraphs be cited in the final answer for transparency. If no paragraph can answer the question, then the system is instructed to output ‘No, I cannot answer’. To further dissect the instructions:
@@
The phrase ‘extracted from the documentation to be assessed’ is a key instruction aimed at clarifying the scope of where the answer should come from. This instruction is particularly necessary for cases where the question might have an implicit answer within the given documentation, but not an explicit one. In other words, it helps ChatGPT make an inference based on the information that is available, rather than stating ‘I cannot answer’ simply because an explicit answer is not given. Once this filtration and aggregation process is complete, we apply the DoX metric to assess the explanatory depth of these refined answers. The culmination of this process is the explanatory relevance score, which is calculated using the formula:
@@ 6. Results
As suggested by the qualitative analysis presented in Section 4, platforms may exhibit varying degrees of compliance regarding ranking transparency. To verify this hypothesis, we went through a manual assessment of the six selected platforms’ documentation, seeking to understand how well they comply with the level of detail and completeness of information required by the P2B Regulation. Then, we compared the results from the manual assessment with those obtained by using the automated assessment tools presented in Section 5. For our manual evaluation, we employed two distinct methods. The first approach involved the expertise of three legal professionals who reviewed all relevant documentation and checklist inquiries. Conversely, our second method engaged over 130 participants. Given the associated costs, this method was applied to just two platforms and was narrowed to only four key questions. The data mentioned in this paper and the code used for this experiment are available at (Sovrano, 2023).
@@ 6.1. Comprehensive Review by Legal Experts
Three Belgian senior legal experts each independently rated the documentation of the six platforms identified using the checklist from Section 4.2. Among these experts, two are academic researchers specializing in EU law on information and communication technologies, with respectively 7 and 5 years of experience in this field. The third, also well-versed in the same domain, with 6 years of experience, is a co-author of our study. To avoid biased results from this legal expert, as well as from the other two, three mitigation measures were taken. First, the set of questions of the checklist was elaborated objectively, before the evaluation of platforms’ documentation. Second, the scores attributed to platforms’ documentation during the assessments were based on an objective scale, described below. Third, legal experts compared and discussed their individual assessments after they were collected, to verify whether bias or other significant discrepancies appeared in the results of the co-author of the study, which was not the case. These experts were instructed to answer the checklist exclusively based on the link sets identified in Section 4.1. Table 1 provides more details on the size of these link sets. The experts were requested to use a scale from 1 to 5, defined as follows:
@@
This scoring system allows for easier averaging across both experts and platforms, providing an aggregate measure of compliance. During this first manual assessment, we rapidly found out that substantial differences existed between several platform operators. As illustrated in Table 2, Bing led with an average compliance score of 3.5. It was followed in descending order by Tripadvisor, Amazon, Google, Booking, and Yahoo. [^6] The low score for Yahoo is attributed to its reliance on Bing’s documentation, as our assessment focused solely on Yahoo’s own documents. Given that a score of ‘2’ indicates a vague or indirect response, our findings resonate with the observations made by the European Commission about the quality of these explanations (Section 2). Yahoo registered the smallest average variance in scores (0.12), indicating minimal disagreement among reviewers. Conversely, Bing and Google exhibited the most variance. As per Table 1, Yahoo and Bing/Google respectively had the fewest and most links under consideration, one might deduce that the volume of documentation influences expert ratings. This makes intuitive sense, as processing and retaining information from extensive texts can be challenging.

### 23 Jan 2024 v2

@@ Abstract.
Compliance with the European Union’s Platform-to-Business (P2B) Regulation helps fostering a fair, ethical and secure online environment. However, it is challenging for online platforms, and assessing their compliance can be difficult for public authorities. This is partly due to the lack of automated tools for assessing the information (e.g., software documentation) platforms provide concerning ranking transparency. Our study tackles this issue in two ways. First, we empirically evaluate the compliance of six major platforms (Amazon, Bing, Booking, Google, Tripadvisor, and Yahoo), revealing substantial differences in their documentation. Second, we introduce and test automated compliance assessment tools based on ChatGPT and information retrieval technology. These tools are evaluated against human judgments, showing promising results as reliable proxies for compliance assessments. Our findings could help enhance regulatory compliance and align with the United Nations Sustainable Development Goal 10.3, which seeks to reduce inequality, including business disparities, on these platforms.
Data and materials: https://doi.org/10.5281/zenodo.10478546.
@@
journalyear: 2024copyright: rightsretainedconference: Software Engineering in Society; April 14–20, 2024; Lisbon, Portugalbooktitle: Software Engineering in Society (ICSE-SEIS’24), April 14–20, 2024, Lisbon, Portugaldoi: 10.1145/3639475.3640112isbn: 979-8-4007-0499-4/24/04ccs: Software and its engineering Documentationccs: Information systems Data analyticsccs: Applied computing Lawccs: Computing methodologies Natural language processing
@@
In the digital world, software powers online platforms such as marketplaces and search engines. For instance, software determines the display order of products on Amazon or hotels on Booking.com. Understanding these systems is vital for businesses, yet the explanations can be complex. The European Union mandates online platforms to reveal the main parameters of their ranking algorithms, aiming for fairness and reducing inequalities. However, ensuring these platforms comply is difficult due to the lack of a standard evaluation method. Our study examined explanations from Google, Yahoo, Bing, Amazon, Booking, and Tripadvisor. We gathered insights from three legal experts and over a hundred people to gauge the clarity of these explanations, finding significant variations in clarity across platforms. Reviewing these explanations manually is cumbersome. To streamline this, we developed automated tools and compared their effectiveness against human assessments. Our discussion revolves around how these tools can enhance transparency on online platforms.
@@ 1. Introduction
To address this issue, the European Union has introduced legislation such as the Platform-to-Business (P2B) Regulation (EU) 2019/1150 (European Parliament and Council, 2019). Among other requirements, this Regulation imposes on platforms to provide easily accessible, comprehensible, and detailed documentation explaining how ranking mechanisms work and affect sellers, including business and corporate website users (Article 5 and Recital 25 (Commission, 2023)). Such a measure aims to create a more equitable environment for businesses that rely on online platforms to connect with consumers. This resonates with Target 10.3 of the United Nations Sustainable Development Goals (of Economic and Affairs, 2023), which aims to ensure equal opportunity and reduce inequalities within and among countries.
However, producing documentation adhering to the P2B Regulation is challenging for online platforms, as recently pointed out by the European Commission (EC) (Commission, 2023). Additionally, monitoring the quality of documentation is challenging for platforms and regulators, affecting compliance assessment, law effectiveness, and social well-being. This is largely due to the lack of standardized, automated tools for assessing compliance across different platform providers. Hence, given the absence of standardized assessment tools, we hypothesize to encounter varying levels of compliance among platforms regarding ranking transparency.
@@
    Source code for the automated tools and referenced data, including manual assessments and checklist (Sovrano, 2023).
@@ 2. Related Work
Additionally, Lukovic (Lukovic, 2021) examined the context and actors of platform rankings. This study details the ranking transparency obligations set by the P2B Regulation and offers critical insights into the framework’s limitations. Bibal et al. (2021) situate the P2B Regulation’s ranking parameters within a broader EU context that mandates explanations for AI systems. This work sheds light on the intentions of EU legislators and specifies the type and extent of information essential for technical compliance. The approaches of both Lukovic (2021) and Bibal et al. (2021) differ from ours as they neither evaluate platforms’ adherence to the Regulation nor introduce tools to enhance regulatory compliance. These elements are actually absent from existing literature, to our knowledge.
@@ 3. Background
This section explores the EU’s P2B Regulation, fostering fairness in online platform-business interactions, alongside the European Commission’s ranking transparency guidelines. It also introduces the DoX metric for measuring explainability, setting the stage for later discussions on assessment tools.
3.1. P2B Regulation & Guidelines
@@
The EU legislator seeks to ensure online platforms offer explanations for business and corporate website users about ranking mechanisms (Commission, 2023; Shaping Europe’s digital future, 2023). This includes understanding how rankings consider offer characteristics, relevance to consumers, and for search engines, website design traits (Art. 5.5). Explanations must also clarify the ability to influence ranking through payments and their impacts (Art. 5.3).
@@
Regarding the form that the explanations should take, both the Regulation and the Guidelines use the term ’drafted’ in English, which refers to written content. [^1] For definitions and examples of the transitive verb form of ”draft,” refer to the Cambridge Dictionary at https://dictionary.cambridge.org/fr/dictionnaire/anglais/draft. To verify this, we looked into the French versions of both texts. The French translation of the Regulation refers to redacted explanations (‘rédigées’) within their terms and conditions for online intermediation services (Art. 3.1 and 5.1). For search engines, the exact text uses the words ‘to state’ (‘énoncée’), which do not imply the use of one form over another (Art. 5.2). The Guidelines, however, refer at some point to the ‘redaction’ of the explanations for both types of providers, potentially implying the use of a written medium in all cases (i.e., “lorsqu’ils rédigent la description […]”, pt. 99). We also looked into the Dutch version of the Regulation, which refers to ‘redacted’ content (‘opgesteld’) regardless of the type of service provider (Art. 3.1 and 5.2).
@@ 3.2. How to Measure Explainability
To calculate DoX scores, the relevance of a text snippet is assessed and aggregated based on its relation to a collection of archetypal questions informed by linguistic theories (see (Sovrano and Vitali, 2023)). Specifically, these relevance measurements are obtained employing pre-trained deep language models tailored for general-purpose answer retrieval (Karpukhin et al., 2020; Bowman et al., 2015). For examples on how DoX works, see the paper (Sovrano and Vitali, 2023).
@@ 4.1. The Study
We selected these platforms for their industry representativeness and the audience profile they serve, which is very broad and allowed us to perform a large-scale manual assessment (cf. Section 6.2).
@@ 4.2. The Checklist
To assess the documentation quality of selected platforms, we created a checklist of essential questions, all needing positive responses for an operator to meet ranking transparency requirements. These questions were crafted by a legal expert and co-author of this study, an academic researcher with expertise in EU law on information and communication technologies and 6 years of experience. This was done using an inductive coding approach (Fereday and Muir-Cochrane, 2006), allowing the checklist questions to emerge from the P2B regulation and guidelines by the European Commission. The checklist was developed based on these materials through the following steps.
@@
However, the Regulation differentiates between two types of platform operators: online intermediation services and online search engines. Both are covered, but they face different requirements, as in Articles 5.1 and 5.2, and point 7 of the Guidelines. Therefore, we have created a different set of yes-or-no questions for each platform type. Importantly, this task was led by solely paying attention to the explicit requirements contained in the Regulation.
Thirdly, we drafted additional questions based on the content of the European Commission’s guidelines. Whenever these guidelines explicitly raised new questions, proposed nuances to the requirements of the legal text, or gave best practices to comply with the P2B Regulation and its objectives, we designed new questions to take these elements into account and incorporated them into our checklist. These questions relate to elements like i) the explanation of what is most important in determining ranking (cf. Q12) and ii) the internal process conducted by the platforms to determine the main parameters (cf. Q14).
Fourthly, we revised certain questions’ wording, without altering their meaning, to address issues in automated answer retrieval from platform documentation. For example, Q11 initially included ‘all the main parameters’ (cf. Guidelines, pt. 24). However, this phrasing led to issues in automated scoring, as no platform explicitly states that the provided list of parameters is exhaustive. To address this, we modified Q11 by omitting the word ’all’, thereby improving the precision of automated assessments.
@@ 2. Q19
Although some of the questions, within each subset, might seem quite similar at first sight, they all call for (at least slightly) different answers, as they contain nuances that originate from the European Commission’s Guidelines. For instance, questions number 4 and 13 appear to be very similar. However, question 4 puts emphasis on the contrast between the parameters selected as main ones and other parameters, whereas question 13 does not. Similarly, questions number 3 and 11 seem quite close to each other. Yet, question 3 (based on the binding text of the Regulation) only refers to the main parameters used, while question 11 (based on the non-binding guidelines) additionally considers how the ranking mechanism works in general.
@@ 5.2. Assessing Explanation Quality with DoX
In particular, we employ the same neural retrieval model (Sovrano et al., 2020) used by the DoX algorithm (cf. Section 3.2) to extract the answers to the checklist questions from the documents and ChatGPT to refine them into comprehensive explanations. Then, DoX is used to measure the explanatory quality of the answers to the checklist questions. Using the earlier example for clarity: while both answers ‘Rankings are based on algorithms’ and ‘Rankings are determined by algorithms that consider factors like user engagement, relevance, and content quality’ would achieve similar relevance ratings, the first would score lower on the DoX metric, indicating its limited explanatory depth.
@@
Paragraphs: {contents} This template is designed for re-elaborating the output of an automated answer retrieval system like the one we employed. The prompt consists of guidelines on how to form a comprehensive answer based on the paragraphs provided for a specific question. It asks the system to generate an answer solely based on the information found in these paragraphs. It also requests that the source paragraphs be cited in the final answer for transparency. If no paragraph can answer the question, then the system is instructed to output ‘No, I cannot answer’.
To further dissect the instructions:
@@
The phrase ‘extracted from the documentation to be assessed’ is a key instruction aimed at clarifying the scope of where the answer should come from. This instruction is particularly necessary for cases where the question might have an implicit answer within the given documentation, but not an explicit one. In other words, it helps ChatGPT make an inference based on the information that is available, rather than stating ‘I cannot answer’ simply because an explicit answer is not given.
Once this filtration and aggregation process is complete, we apply the DoX metric to assess the explanatory depth of these refined answers.
The culmination of this process is the explanatory relevance score, which is calculated using the formula:
@@ 6. Results
As suggested by the qualitative analysis presented in Section 4, platforms may exhibit varying degrees of compliance regarding ranking transparency. To verify this hypothesis, we went through a manual assessment of the six selected platforms’ documentation, seeking to understand how well they comply with the level of detail and completeness of information required by the P2B Regulation. Then, we compared the results from the manual assessment with those obtained by using the automated assessment tools presented in Section 5.
For our manual evaluation, we employed two distinct methods. The first approach involved the expertise of three legal professionals who reviewed all relevant documentation and checklist inquiries. Conversely, our second method engaged over 130 participants. Given the associated costs, this method was applied to just two platforms and was narrowed to only four key questions.
The data mentioned in this paper and the code used for this experiment are available at (Sovrano, 2023).
@@ 6.1. Comprehensive Review by Legal Experts
Three Belgian senior legal experts each independently rated the documentation of the six platforms identified using the checklist from Section 4.2. Among these experts, two are academic researchers specializing in EU law on information and communication technologies, with respectively 7 and 5 years of experience in this field. The third, also well-versed in the same domain, with 6 years of experience, is a co-author of our study. To avoid biased results from this legal expert, as well as from the other two, three mitigation measures were taken. First, the set of questions of the checklist was elaborated objectively, before the evaluation of platforms’ documentation. Second, the scores attributed to platforms’ documentation during the assessments were based on an objective scale, described below. Third, legal experts compared and discussed their individual assessments after they were collected, to verify whether bias or other significant discrepancies appeared in the results of the co-author of the study, which was not the case.
These experts were instructed to answer the checklist exclusively based on the link sets identified in Section 4.1. Table 1 provides more details on the size of these link sets.
The experts were requested to use a scale from 1 to 5, defined as follows:
@@
This scoring system allows for easier averaging across both experts and platforms, providing an aggregate measure of compliance.
During this first manual assessment, we rapidly found out that substantial differences existed between several platform operators. As illustrated in Table 2, Bing led with an average compliance score of 3.5. It was followed in descending order by Tripadvisor, Amazon, Google, Booking, and Yahoo. [^5] The low score for Yahoo is attributed to its reliance on Bing’s documentation, as our assessment focused solely on Yahoo’s own documents. Given that a score of ‘2’ indicates a vague or indirect response, our findings resonate with the observations made by the European Commission about the quality of these explanations (Section 2).
Yahoo registered the smallest average variance in scores (0.12), indicating minimal disagreement among reviewers. Conversely, Bing and Google exhibited the most variance. As per Table 1, Yahoo and Bing/Google respectively had the fewest and most links under consideration, one might deduce that the volume of documentation influences expert ratings. This makes intuitive sense, as processing and retaining information from extensive texts can be challenging.